A company that operates for 18 hours a day has introduced a new daily wage system that
incorporates variable hourly rate to encourage longer shifts. The company pays $10 for first
hour or any part of the first hour. Afterwards, company pays additional 2% per hour for each
hour or part of each additional hour. Write a piece-wise function that shows the salary as a
function of hours. Also sketch a graph of this function.
1
Expert's answer
2021-09-24T12:22:54-0400

Let "f(x)" shows the total salary on hourly basis. ""x"" denotes the number of hours.

"f(x) = \begin{cases}
 10 &\text{if } x<=1 \\
 (0.02+10)x &\text{if } 1<x<=18
\end{cases}"

Simplifying "f(x)" , we get:

"f(x) = \begin{cases}
 10 &\text{if } x<=1 \\
 (10.02)x &\text{if } 1<x<=18
\end{cases}"


Graph of "f(x)" will be: